Grade 11Mechanics

A force of 6 N and another of 8 N can be applied to produce the effect of a single force equal to

Sushanth mahindra
9 Years agoGrade 11

1 Answer

Agastya Varma
9 Years ago
 
The resultant of any two forces depends upon the angle at which they are inclined to each othe. The maximum resultant is produced when the vectors act in the the same direction i.e. 0and minimum when the act in opposite directions i.e. 180o. So here the max. Force is 6+8= 14N and min. 6-8= 2N. 
Hpoe this helped you. 
The resultant of any two forces depends upon the angle at which they are inclined to each othe. The maximum resultant is produced when the vectors act in the the same direction i.e. 0and minimum when the act in opposite directions i.e. 180o. So here the max. Force is 6+8= 14N and min. 6-8= 2N. 
Hpoe this helped you.